# Advanced Weights for IXPE Polarization Analysis

### v1

@@ Abstract
As the Imaging X-ray Polarimetry Explorer (IXPE) measures increasingly faint sources, the need for precise polarimetry extraction becomes paramount. In addition to previously described neural-net (NN) weights, we introduce here point-spread function weights and particle background weights, which can be critical for faint sources. In some cases these can be augmented by time/phase and energy weights. We provide a publicly available analysis tool to incorporate these new weights, validate our method on simulated data, and test it on archival IXPE observations. Together these weights decrease the area of the polarization uncertainty contour by a factor of two and will be essential for background-limited IXPE observations.
@@ 1 Introduction
The Imaging X-ray Polarimetry Explorer (IXPE, Weisskopf et al., 2022; Soffitta et al., 2021) provides unique insights into the geometry of active galactic nuclei, X-ray binaries, pulsar wind nebulae (PWNe) and other X-ray sources. This advance comes from the track images which IXPE’s gas pixel detectors (GPDs, Baldini et al., 2021) provide for each measured X-ray. Much has been achieved with the analytic mission-standard moments (Mom) track analysis which provides estimates of the events’ position, energy, electric vector polarization angle (EVPA), and EVPA accuracy. As we observe fainter IXPE targets exposures become very long and background contamination becomes highly significant; advanced methods to more optimally exploit the track image information become increasingly attractive.
For example Peirson et al. (2021) and Cibrario et al. (2025) introduce neural net (NN)-based methods to extract event parameters from IXPE tracks. Since photon and particle event morphologies differ, Di Marco et al. (2023) suggest cuts based on track metadata that partly suppress particle background. These measurements best constrain source polarization when maximum likelihood estimators (MLE Vianello et al., 2015; Marshall, 2021; González-Caniulef et al., 2023; Marshall, 2024) are used. Additional refinements are possible. Prior X-ray observations generally provide detailed image morphology, spectrum, and temporal variation measurements of IXPE targets (e.g. Wong et al., 2023). With accurate IXPE response functions, these measurements allow IXPE event positions, energies, and arrival times to be used as weights, concentrating the statistical power of IXPE data onto the polarization properties.
In this work, we develop a system of track-based weights to improve extraction of source polarization. The NN-technique of (Peirson et al., 2021) already provides effective polarization weights. We introduce a new NN which provides weights for particle rejection. We use new, accurate on-orbit IXPE PSFs (deconvolved from point source observations; Dinsmore & Romani, 2024) to provide spatial weights. These weights correct for “polarization leakage” (Bucciantini et al., 2023a), spurious polarized fringes introduced by correlations in errors between the reconstructed event position and EVPA (predicted by Dinsmore & Romani, 2024). Section §2 details these weights and their incorporation into an MLE analysis. §3 uses simulations of a polarized point source to show large improvement over the mission-standard extraction, and insensitivity to systematics. We apply the method to published IXPE data in §4, again demonstrating improved uncertainties, and conclude with a brief discussion in §5.
@@ 2 Methods
For event $i$, $P_{s}(\psi_{i})$ and $P_{s}(\bm{r}_{i}|\psi_{i})$ are the PDFs of EVPA $\psi_{i}$ and position $\bm{r}_{i}$, assuming an origin from source $s$, while $p_{i,s}$ is the probability that $i$ was emitted by $s$. We use “source” to refer to any potential origin of an IXPE event, which may include astrophysical sources, photon background, or particle background. In Eq. 1, $P_{s}(\psi_{i})$ and $P_{s}(\bm{r}_{i}|\psi_{i})$ act as polarization (§2.1) and spatial weights (§2.2). $p_{i,s}$ may be used to de-weight particle-like events (§2.3) and incorporate spectral and temporal weights (§2.4).
The maximum likelihood estimator (MLE) of parameter $\theta$ is the value that maximizes $L$. With many events, MLE methods are unbiased and deliver optimal uncertainties. Here we estimate the covariance matrix $\Sigma$ by computing $(\Sigma^{-1})_{ij}=-\frac{\partial}{\partial\theta_{i}}\frac{\partial}{\partial\theta_{j}}\ln L$ at the maximum. The Bayesian approach with wide, uniform priors is equivalent if Gaussian posteriors are assumed, though this assumption may be lifted using a Markov-chain Monte Carlo (MCMC) method. Tools for both types of fits are provided as updates to the LeakageLib package, originally designed to treat polarization leakage (Dinsmore & Romani, 2024). [^1] https://github.com/jtdinsmore/leakagelib
@@ 2.1 Polarization Weights
We define $q_{i}=\cos 2\psi_{i}$ and $u_{i}=\sin 2\psi_{i}$ as the Stokes coefficients of event $i$. The PDF of $\psi_{i}$ is commonly modeled as
@@ 2.2 Spatial Weights
where $I_{s}(\bm{x})$ is the (generally known) unpolarized surface brightness map of the source, $P_{\mathrm{mir}}$ is the mirror PSF, and $\bm{\delta}$ is the reconstruction error vector. The reconstruction error PDF $P(\bm{\delta}|\psi_{i})$ depends on the event EVPA, leading to polarization leakage. As in Bucciantini et al. (2023a), we Taylor expand $P_{\mathrm{mir}}$ in $\bm{\delta}$ since $\delta\ll$ the PSF width. The $\bm{\delta}$ integral in Eq. 4 reduces to the first few moments of $P(\bm{\delta}|\psi_{i})$, whose values and energy dependence were measured in Dinsmore & Romani (2024). In particular,
@@
where $I_{0}$, $Q_{0}$, $U_{0}$, $X_{0}$, and $Y_{0}$ are the known functions of position and event energy reported in that work, and
$C_{i,s}$ $=\int d^{2}\bm{r}\int d^{2}\bm{x}\,I_{s}(\bm{x})\bigg{[}I_{0}(\bm{r}-\bm{x})$ (6)
$+\frac{\mu_{i}}{2}\left(QQ_{0}(\bm{r}-\bm{x})+UU_{0}(\bm{r}-\bm{x})\right)\bigg{]}$
is the normalization constant, obtained by integrating $P_{s}(\bm{r}_{i}|\psi_{i})P(\psi_{i})$ over $\psi_{i}$ and $\bm{r}_{i}$.
Spatial weights are important for observations of faint sources where where they help to de-weight background events. They also correct for polarization leakage biases, as we now prove. Consider polarimetry extraction of a bright point source with a small or irregular region, where polarization leakage biases the naive estimator Eq. 3. When spatial weights are included, the likelihood reduces to $L=\prod_{i}P(\psi_{i})C_{i,s}^{-1}$ (terms that do not depend on the fit parameters are dropped), and the MLE is
$\hat{Q}_{\mathrm{src,\,spatial}}=\frac{\sum\mu_{i}[2q_{i}-q_{0}(\bm{r}_{i})]}{\sum\mu_{i}^{2}},$ (7)
$\hat{U}_{\mathrm{src,\,spatial}}=\frac{\sum\mu_{i}[2u_{i}-u_{0}(\bm{r}_{i})]}{\sum\mu_{i}^{2}}$
where $q_{0}(\bm{r})=Q_{0}(\bm{r})/I_{0}(\bm{r})$ and $u_{0}(\bm{r})=U_{0}(\bm{r})/I_{0}(\bm{r})$ are the expected leakage patterns. Eq. 7 explicitly demonstrates that our MLE technique correctly subtracts off polarization leakage effects in this simplified limit. We argue that this guarantee extends to more complicated scenarios, provided large data sets and valid source flux models, due to the accuracy of MLE methods.
@@ 2.3 Background Weights
The CNN is trained with binary cross-entropy loss on data sets extracted from IXPE observations of bright, low-polarization point sources (observation IDs 01002401, 01002601, 02002399, and 02001901). The source $S$ data set consists of events within $26^{\prime\prime}$ of the source, while the equal-size background $B$ data set contains events beyond $117^{\prime\prime}$. We take equal sampling of the 1-10 keV energy range in both data sets so that the trained CNN will not exhibit energy-dependent bias. But the background $B$ is inevitably polluted by wings of the central PS and astrophysical photon backgrounds. Training on $B$ and $S$ gives a initial CNN, used to create a nearly photon-free $B^{\prime}$ by removing low $\pi_{i}$ (photon) events from $B$. We then retrain on $S$ and $B^{\prime}$. The re-trained CNN performs quite well; $95\%$ of validation events in the $B^{\prime}$ (photon-cleaned) data set are cut with a $\pi_{i}<0.5$ threshold. Only 5% of events were misclassified, assuming $B^{\prime}$ is perfectly clean. This final CNN and a script to tag IXPE level 2 files with $\pi_{i}$ is provided in the updated version of LeakageLib.
@@
Fig. 1 demonstrates the efficacy of particle weights for faint sources using the brightest of all time (B.O.A.T.) Gamma-ray burst as an example. GRB 221009A triggered an IXPE target of opportunity observation (Negro et al., 2023) that began over two days after the burst’s detection and continued for two more. Echos of the initial explosion were detected as faint rings surrounding the point-like afterglow, reflected by Galactic dust. These rings are much more prominent when particles are de-weighted rather than cut, which will be critical for future polarization analyses. Good agreement between the model GRB radial profile (see appendix A) and the IXPE radial profile further illustrates this.
Our CNN assigns high $\pi_{i}$ to events cut by Di Marco et al. (2023). Fig. 2 shows our particle character estimates in the space of two level 1 parameters used to determine these cuts. Contours show the event distribution within source ($r<10^{\prime\prime}$) and background regions ($180<r<280^{\prime\prime}$) for the IXPE observations of PSR B0540-69. Low $\pi_{i}$ events are removed from the background contour to clean it of solar photons. Clearly, the additional information of the NN allows improved background suppression, since some uncut events are flagged with high $\pi_{i}$. De-weighting these events may incur slight source flux loss; our treatment takes this into account when evaluating PD.
@@ 2.4 Spectral and Temporal Weights
MLE methods can suffer from bias when applied to observations of multiple bright sources with distinct spectra. In such cases, the events from each source exhibit different modulation factor distributions and the correct distribution must be used to extract polarizations for a given source. For example, consider a polarization MLE for a point source embedded in an unpolarized background. A naive approach would estimate aperture polarization using Eq. 3, measure the counts $N_{\mathrm{src}}$ ($N_{\mathrm{bg}}$) in source (background) apertures, and multiply the estimator by the background rate $f_{N}=N_{\mathrm{src}}/(N_{\mathrm{src}}-\phi N_{\mathrm{bg}})$ where $\phi$ is the source-to-background area ratio. This method is incorrect because Eq. 3 divides by $\mu^{2}$ summed over the entire aperture, which includes both source and spectrum-polluting background events. An unbiased approach is to multiply by a modified background rate
@@
For more general MLE methods, we account for spectral differences by replacing the flux $F_{s}$ in Eq. 9 with a spectrum $F_{s}(E_{i})$, multiplying by instrument response functions to account for effective area and energy resolution. The spectral parameters can in principle be simultaneously fit with the polarization to achieve spectropolarimetric fits—potentially useful for detailed spectral studies of brighter sources. But spectral fitting requires modeling of IXPE’s finite energy resolution and is beyond the scope of this work, which focuses on broadband polarimetry of faint sources.
Temporal or phase weights may also be included by replacing $F_{s}$ with the lightcurve $F_{s}(t_{i})$. Wong et al. (2023) explored this capability in a “simultaneous fitting” analysis of the Crab pulsar and PWN.
@@ 3 Validation of Spatial and Spectral Weighting on Simulated Data
For very bright (low background) sources, spatial weighting presents a modest improvement in precision over PCUBE. This improvement is largely due to the bright-source MLE method (Eq. 3), which outperforms the PCUBE approach (Eq. 11). As the background increases, PCUBE uncertainties (and uncertainties from any other aperture-based method) grow, even after background subtraction. However, the spatial weighting strongly suppresses this background so that uncertainties increase very slowly. Thus for faint (especially extended) sources, spatial weighting provides a dramatic improvement.

### v2

@@ Abstract
As the Imaging X-ray Polarimetry Explorer (IXPE) measures increasingly faint sources, the need for precise polarimetry extraction becomes paramount. In addition to previously described neural-net (NN) weights, we introduce here point-spread function weights and particle background weights, which can be critical for faint sources. In some cases these can be augmented by time/phase and energy weights. We provide a publicly available analysis tool to incorporate these new weights, validate our method on simulated data, and test it on archival IXPE observations. Together these weights decrease the area of the polarization uncertainty contour by a factor of two compared to baseline IXPE analysis and will be essential for background-limited IXPE observations.
@@ 1 Introduction
The Imaging X-ray Polarimetry Explorer (IXPE, Weisskopf et al., 2022; Soffitta et al., 2021) provides unique insights into the geometry of active galactic nuclei, X-ray binaries, pulsar wind nebulae (PWNe) and other X-ray sources. This advance comes from the track images which IXPE’s gas pixel detectors (GPDs, Baldini et al., 2021) provide for each measured X-ray. Much has been achieved with the analytic mission-standard moments (Mom) track analysis which provides estimates of the events’ position, energy, electric vector polarization angle (EVPA), and EVPA accuracy. As we observe fainter IXPE targets, exposures become very long and background contamination becomes highly significant; advanced methods to more optimally exploit the track image information become increasingly attractive.
For example Peirson et al. (2021) and Cibrario et al. (2025) introduce neural net (NN)-based methods to extract event parameters from IXPE tracks. Since photon and particle event morphologies differ, Di Marco et al. (2023) suggest cuts based on track metadata that partly suppress particle background. Spectro-polarimetric fitting uses event energies to improve polarization analysis (e.g. 3ML, Vianello et al., 2015), and IXPE’s excellent timing precision may be used to separate a variable source from a constant background (e.g. the ‘simultaneous fitting’ of Wong et al., 2023). Measurements best constrain source polarization when maximum likelihood estimators (MLE, Marshall, 2021; González-Caniulef et al., 2023; Marshall, 2024) are applied. Though useful, many of these methods are not compatible, as presently formulated. Furthermore, additional refinements are possible. Prior X-ray observations generally provide detailed image morphology of IXPE targets. With accurate IXPE response functions, these measurements allow IXPE event positions and energies to be used as weights, concentrating the statistical power of IXPE data onto the polarization properties.
In this work, we develop a system of track-based weights to improve extraction of source polarization. The NN-technique of (Peirson et al., 2021) already provides effective polarization weights. We introduce a new NN which provides weights for particle rejection. We also use new, accurate on-orbit IXPE PSFs (deconvolved from point source observations; Dinsmore & Romani, 2024) to provide spatial weights. These weights correct for “polarization leakage” (Bucciantini et al., 2023a), spurious polarized fringes introduced by correlations in errors between the reconstructed event position and EVPA (as described by Dinsmore & Romani, 2024). §2 details these and temporal/phase and energy weights, and presents their incorporation into a unified MLE analysis. §3 uses simulations of a polarized point source to show large improvement over the mission-standard extraction, and insensitivity to systematics. We apply the method to published IXPE data in §4, again demonstrating improved uncertainties, and conclude with a brief discussion in §5.
@@ 2 Methods
For event $i$, $P_{s}(\psi_{i})$ and $P_{s}(\bm{r}_{i}|\psi_{i})$ are the PDFs of EVPA $\psi_{i}$ and position on the sky $\bm{r}_{i}$, assuming they originated from source $s$, while $p_{i,s}$ is the probability that $i$ was emitted by $s$. We use “source” to refer to any potential origin of an IXPE event, which may include astrophysical sources, photon background, or particle background. In Eq. 1, $P_{s}(\psi_{i})$ and $P_{s}(\bm{r}_{i}|\psi_{i})$ act as polarization (§2.1) and spatial weights (§2.2). $p_{i,s}$ may be used to de-weight particle-like events (§2.3) and incorporate spectral and temporal weights (§2.4).
The maximum likelihood estimator (MLE) of parameter $\theta$ is the value that maximizes $L$. With many events, MLE methods are unbiased and deliver optimal uncertainties. Here we estimate the covariance matrix $\Sigma$ by computing $(\Sigma^{-1})_{ij}=-\frac{\partial}{\partial\theta_{i}}\frac{\partial}{\partial\theta_{j}}\ln L$ at the maximum. The Bayesian approach with wide, uniform priors in the source Stokes coefficients is equivalent if Gaussian posteriors are assumed, though this assumption may be lifted using a Markov-chain Monte Carlo (MCMC) method. Tools for both types of fits are provided as updates to the LeakageLib package, originally designed to treat polarization leakage (Dinsmore & Romani, 2024). [^1] https://github.com/jtdinsmore/leakagelib
@@ 2.1 Polarization Weights
We define $q_{i}=\cos 2\psi_{i}$ and $u_{i}=\sin 2\psi_{i}$ as the Stokes coefficients of event ii. [^2] Our definition of $q_{i}$ and $u_{i}$ differs from that of the IXPE level 2 files, which report Stokes coefficients as $2\cos 2\psi_{i}$ and $2\sin 2\psi_{i}$. The PDF of $\psi_{i}$ is commonly modeled as
@@ 2.2 Spatial Weights
where $I_{s}(\bm{x})$ is the (usually known) unpolarized surface brightness map of the source, $P_{\mathrm{mir}}$ is the mirror PSF, and $\bm{\delta}$ is the reconstruction error vector. The reconstruction error PDF $P(\bm{\delta}|\psi_{i})$ depends on the event EVPA, leading to polarization leakage. As in Bucciantini et al. (2023a), we Taylor expand $P_{\mathrm{mir}}$ in $\bm{\delta}$ since $\delta\ll$ the PSF width. The $\bm{\delta}$ integral in Eq. 4 reduces to the first few moments of $P(\bm{\delta}|\psi_{i})$, whose values and energy dependence were measured in Dinsmore & Romani (2024). In particular,
@@
where $I_{0}$, $Q_{0}$, $U_{0}$, $X_{0}$, and $Y_{0}$ are the known functions of position and event energy reported in Dinsmore & Romani (2024), and
$C_{i,s}$ $=\int_{A}d^{2}\bm{r}\int d^{2}\bm{x}\,I_{s}(\bm{x})\bigg[I_{0}(\bm{r}-\bm{x})$ (6)
$+\frac{\mu_{i}}{2}\left(Q_{s}Q_{0}(\bm{r}-\bm{x})+U_{s}U_{0}(\bm{r}-\bm{x})\right)\bigg]$
is the normalization constant, with $A$ representing the extraction aperture. Eq. 6 is obtained by setting $\int_{0}^{2\pi}d\psi\int_{A}d^{2}\bm{r}\,P_{s}(\bm{r}|\psi)P_{s}(\psi)=1$.
Spatial weights, represented by the $P_{s}(\bm{r}_{i}|\psi_{i})$ term of Eq. 1, are important for observations of faint sources because they help to de-weight background events. They also correct for polarization leakage biases, as we now prove. Consider polarimetry extraction of a bright point source with a small or irregular aperture, where polarization leakage pollutes the spatially unweighted estimator with additional Stokes coefficients $Q_{\ell}$ and $U_{\ell}$. The methods of Dinsmore & Romani (2024) predict
$Q_{\ell}=\frac{\sum_{i}\mu_{i}\int_{A}d^{2}\bm{r}\,Q_{0}(\bm{r})}{\sum_{i}\mu_{i}^{2}},\ U_{\ell}=\frac{\sum_{i}\mu_{i}\int_{A}d^{2}\bm{r}\,U_{0}(\bm{r})}{\sum_{i}\mu_{i}^{2}}.$ (7)
for a weakly polarized point source.
Since the source is bright, the data may be modeled by a single component $s=\mathrm{src}$ representing the point source, giving $L=\prod_{i}P_{\mathrm{src}}(\psi_{i})P_{\mathrm{src}}(\bm{r}_{i}|\psi_{i})$. Furthermore, $P_{\mathrm{src}}(\bm{r}_{i}|\psi_{i})$ is independent of the fit parameters $Q_{\mathrm{src}}$ and $U_{\mathrm{src}}$ except through $C_{i,\mathrm{src}}^{-1}$, and we may drop the constant terms because they do not affect the MLE. Hence $L\propto\prod_{i}P_{\mathrm{src}}(\psi_{i})C_{i,\mathrm{src}}^{-1}$. For a point source, Eq. 6 gives $C_{i,\mathrm{src}}=1+\frac{\mu_{i}}{2}\int_{A}d^{2}\bm{r}\,\left[Q_{\mathrm{src}}Q_{0}(\bm{r})+U_{\mathrm{src}}U_{0}(\bm{r})\right]$. Maximizing this simplified $L$ in the low PD limit gives the MLE:
$\hat{Q}_{\mathrm{src}}=2\frac{\sum\mu_{i}q_{i}}{\sum\mu_{i}^{2}}-Q_{\ell},\qquad\hat{U}_{\mathrm{src}}=2\frac{\sum\mu_{i}u_{i}}{\sum\mu_{i}^{2}}-U_{\ell}.$ (8)
This demonstrates that our MLE technique correctly subtracts polarization leakage effects. This will extend to more complicated scenarios, provided good statistics and valid source flux models, given the unbiased source parameter retrieval of MLE methods.
@@ 2.3 Background Weights
The CNN is trained with binary cross-entropy loss on data sets extracted from IXPE observations of bright, low-polarization point sources (observation IDs 01002401, 01002601, 02002399, and 02001901). The source $S$ data set consists of events within $26^{\prime\prime}$ of the source, while the equal-size background $B$ data set contains events beyond $117^{\prime\prime}$. We take equal sampling of the 1-10 keV energy range in both data sets so that the trained CNN will not exhibit energy-dependent bias. But the background $B$ is inevitably polluted by wings of the central point source and astrophysical photon backgrounds. Training on $B$ and $S$ gives an initial CNN, used to create a nearly photon-free $B^{\prime}$ by removing low $\pi_{i}$ (photon) events from $B$. We then retrain on $S$ and $B^{\prime}$. The re-trained CNN performs quite well; $95\%$ of validation events in the $B^{\prime}$ (photon-cleaned) data set are cut with a $\pi_{i}<0.5$ threshold. Only 5% of events were misclassified, assuming $B^{\prime}$ is perfectly clean. This final CNN and a script to tag IXPE level 2 files with $\pi_{i}$ is provided in the updated version of LeakageLib.
@@
Fig. 1 demonstrates the efficacy of particle weights for faint sources using the brightest of all time (B.O.A.T.) Gamma-ray burst as an example. GRB 221009A triggered an IXPE target of opportunity observation (Negro et al., 2023) that began just two days after the burst’s detection and continued for two more. Echos of the initial explosion were detected as faint rings surrounding the point-like afterglow, reflected by Galactic dust. These rings are much more prominent when particles are de-weighted rather than cut, which will be critical for future polarization analyses. Good agreement between the model GRB radial profile (see appendix A) and the IXPE radial profile further illustrates this.
Our CNN assigns high $\pi_{i}$ to events cut by Di Marco et al. (2023). Fig. 2 shows our particle character estimates in the space of two level 1 parameters used to determine these cuts. Contours show the event distribution within source ($r<10^{\prime\prime}$) and background regions ($180^{\prime\prime}<r<280^{\prime\prime}$) for the IXPE observations of PSR B0540-69. Low $\pi_{i}$ events are removed from the background contour to clean it of solar photons. Clearly, the additional information of the CNN allows improved background suppression, since some uncut events are flagged with high $\pi_{i}$. De-weighting these events may incur slight source flux loss; our treatment takes this into account when evaluating PD.
@@ 2.4 Spectral and Temporal Weights
MLE methods can suffer from bias when applied to observations of multiple bright sources with distinct spectra. In such cases, the events from each source exhibit different modulation factor distributions and the correct distribution must be used to extract polarizations for a given source. For example, consider a polarization MLE for a point source embedded in an unpolarized background. A naive approach would estimate aperture polarization using Eq. 3, measure the counts $N_{\mathrm{src}}$ ($N_{\mathrm{bg}}$) in source (background) apertures, and multiply the estimator by the background rate $f_{N}=N_{\mathrm{src}}/(N_{\mathrm{src}}-\phi N_{\mathrm{bg}})$ where $\phi$ is the source-to-background area ratio. While suitable for PCUBE, for MLE methods $f_{N}$ is incorrect because Eq. 3 divides by $\mu^{2}$ summed over the entire aperture, which includes both source and spectrum-polluting background events. An unbiased approach is to multiply by a modified background rate
@@
For MLE analyses using the full likelihood instead of Eq. 3, we account for spectral differences by replacing the flux $F_{s}$ in Eq. 10 with a spectrum $F_{s}(E_{i})$, multiplying by instrument response functions to account for effective area and energy resolution. The spectral parameters can in principle be simultaneously fit with the polarization to achieve spectropolarimetric fits—potentially useful for detailed spectral studies of brighter sources. But spectral fitting requires modeling of IXPE’s finite energy resolution and is beyond the scope of this work, which focuses on broadband polarimetry of faint sources.
Wong et al. (2023) explored phase weighting with IXPE in a “simultaneous fitting” analysis of the Crab pulsar and PWN. In the maximum likelihood framework, temporal or phase weights may be included by replacing $F_{s}$ with the light curve $F_{s}(t_{i})$. Phase weighting with this method makes no immediate improvement over Wong et al. (2023), but it allows the simultaneous use of other weight techniques which are beneficial. If the expected source polarization is also time-variable, $Q_{s}$ and $U_{s}$ may likewise be replaced with the variability model $Q_{s}(t_{i})$ and $U_{s}(t_{i})$. This paper considers only constant $Q_{s}$ and $U_{s}$ in a given phase bin, but LeakageLib provides tools to fit for continuously variable source polarizations, as seen in e.g. IXPE observations of pulsars.
@@ 3 Validation of Spatial and Spectral Weighting on Simulated Data
For very bright (low background) sources, spatial weighting presents a modest improvement in precision over PCUBE. This improvement is largely due to the bright-source MLE method (Eq. 3), which outperforms the PCUBE approach (Eq. 12). The next section demonstrates, using real IXPE data, that this improvement over PCUBE occurs even when Moments weights $w_{i}$ are used. As the background increases, PCUBE uncertainties (and uncertainties from any other aperture-based method) grow, even after background subtraction. However, the spatial weighting strongly suppresses this background so that uncertainties increase very slowly. Thus for faint (especially extended) sources, spatial weighting provides a dramatic improvement.